OpenStudy (anonymous):

Model - R=P(i/(1-(1+i)^(-nt))), where i =r/n, r=8% compounded monthly(n) and p=90000, find t. When R=1200 1200 = 90000(0.08/(12)/(1-(1+0.08/(12))^-(12)t Find t(year)

OpenStudy (mathmath333):

this ur model \(\large \color{black}{\begin{align} R=P\left(\dfrac{i}{1-(1+i)^{-nt}} \right)\hspace{.33em}\\~\\ \end{align}}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Its pretty easy to solve I guess i would take the logarithm on both ends \[\large \color{black}{\begin{align} 90000=12000\left(\dfrac{\frac{.08}{12}}{1-(1+\frac{.08}{12})^{-12t}} \right)\hspace{.33em}\\~\\ \end{align}}\] \[90000(1-(1+\frac{.08}{12})^{-12t}=12000(\frac{.08}{12})\] \[(1+\frac{.08}{12})^{-12t}=-12000(\frac{.08}{12})+90000\] Simplify and take the log on both sides

OpenStudy (anonymous):

whoops its 1200 not 12000

OpenStudy (anonymous):

\[(1+\frac{.08}{12})^{-12t}=-1200(\frac{.08}{12})+90000\]

OpenStudy (anonymous):

Why did you add the 90000?

OpenStudy (anonymous):

Its 90,000

OpenStudy (anonymous):

and I just did basic simplifying

OpenStudy (anonymous):

so its not?\[-90000(1+.08/12)^\left( -12t \right)=1200(.08/12)-90000\]

OpenStudy (anonymous):

lol u r actually right -.- Sorry i skipped out that step

OpenStudy (anonymous):

\[(1+.08/12)^\left( -12t \right)=\frac{1200(.08/12)-90000}{-90,000}\]

OpenStudy (anonymous):

\[\log.08/12(.08/12)^\left( -12t \right)=\log.08/12(9)\] is what i have as of now

OpenStudy (anonymous):

hmmmm this what I have \[\log(1+.08/12)^{-12t}=\log(\frac{1200(.08/12)-90000}{-90,000})\]

OpenStudy (anonymous):

forgot the 1 \[\log_{1+.08/12} (1+.08/12)^\left( -12t \right)=\log_{1+.08/12}(9)\] \[-12t=\log_{1+.08/12}(9)\]

OpenStudy (anonymous):

I dont get why u just dont use log base 10 ... easier to work with

OpenStudy (anonymous):

I would but, I want to cancel the 1+... so t can get separated

OpenStudy (anonymous):

Or atleast thats what I was taught

OpenStudy (anonymous):

Hmmm this is how I would go about it tbh Why cant you just add 1+.08/12 Do you have a scientific calculator?

OpenStudy (anonymous):

Actuallly do what you were taught in school :P

OpenStudy (anonymous):

Just not sure how you got the right hand side of the equation

OpenStudy (anonymous):

I get \[-12t=330.68...\] \[t=-27.556...\]

OpenStudy (anonymous):

I got 9 from solving it on a calculator

OpenStudy (anonymous):

but, I don't think t being a negative answer, would be "correct" - in that if there is a negative year

OpenStudy (anonymous):

Ya something doesnt look right there Wait lemme go over ur steps

OpenStudy (anonymous):

How did you get the 9 again?

OpenStudy (anonymous):

Cuz I dont seem to get that answer

OpenStudy (anonymous):

I got 9 by doing \[(1200(.08)-90000)/(-90000)\]

OpenStudy (anonymous):

for the 12, but it was included when typed into the calculator

OpenStudy (anonymous):

im giving up .... @precal most probably will catch my error

OpenStudy (anonymous):

Well, I was at wrong - forgot the parenthesis

OpenStudy (anonymous):

0.9999111111 is what you get - you forgot it was divided by a negative 90k

OpenStudy (anonymous):

ohhhhh yup

OpenStudy (anonymous):

It shoudl work now though

OpenStudy (precal):

Good job!

OpenStudy (anonymous):

\[\log{1+.08/12} ^\left( -12t \right)=\log{0.9999111111} \] \[0.001806^\left( -12t \right)=\log(0.9999111111)\]

OpenStudy (anonymous):

this would mean that its too complicated and I should just do log1+.08/12(1+.08/12)

OpenStudy (anonymous):

\[-12t=\log{1+.08/12}(0.9999111111)\]

OpenStudy (anonymous):

hmmm do you mean \[-12t=\log_{1+.08/12}(.999911111)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I keep getting a really small number for t Makes no difference how I compute this -.-

OpenStudy (anonymous):

same, I get 0.001115

OpenStudy (anonymous):

\[\log_{1+(.08/12)}(1+.08/12)^{-12t}=\log_{1+(.08/12)}(\frac{1200(.08/12)-90000}{-90,000})\] \[-12t=\log_{1+(.08/12)}(\frac{1200(.08/12)-90000}{-90,000})\] @dan815 Can you solve for t I keep getting an abnormally small number

OpenStudy (anonymous):

mmk I got it to where t~8.693 by doing \[1200=90000(i/1-(1+i)^\left( -12t \right)\] \[-12t=\log_{1+.08/12} (.5)\]

OpenStudy (anonymous):

From what I remember the original question was an couple owes $90000, where their i is at a rate of 8% compounded monthly. And they paid $1200 per month - how many years would it take for them to pay it off

OpenStudy (anonymous):

I got 8.69 using TVM Solver

OpenStudy (anonymous):

So that is def the correct answer

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (anonymous):

How did you get the .5 though?

OpenStudy (anonymous):

i did \[600-1200/(-1200)\] where 600 came from 90000(i)